Suppose you have $36,000 to invest. You’re considering Miller-Moore Equine Enterprises (MMEE), which is currently selling for $80 per share. You notice that a put option with a $80 strike is available with a premium of $3.00. Calculate your percentage return on the put option for the six-month holding period if the stock price declines to $76 per share. (A negative value should be indicated by a minus sign. Leave no cells blank - be certain to enter "0" wherever required. Do not round intermediate calculations. Enter your 6-month return as a percent rounded to 2 decimal places.)

percentage return?
Answer #1

Amount available to invest = $36,000

premium = $3.00

so number of shares that can be purchased = 36,000 / 3 = 12,000

value of put option = (strike price - exercise price) * number of shares

= (80 -76) * 12,000

= 48,000

so 6 months return = (48,000 - 36,000) / 36,000

= 12,000 / 36,000

= 0.33333 or 33.33%(rounded to two decimal places)